# Undersmoothed LASSO Models for Propensity Score Weighting and Synthetic Negative Control Exposures for Bias Detection

### v2

@@ 2.4 Using Synthetic Data for Bias Detection with LASSO PS Analyses
Because the synthetic cohorts do not approximate the full data distribution, they are simply used to detect analyses that are unlikely to satisfy partial exchangeability (discussed in Section 2.4.1). Using synthetic cohorts for bias detection rather than model selection helps to mitigate the challenge of making analytic decisions that overfit to the synthetic data. Below, we formally define negative control exposures and describe how they relate to conditional exchangeability in the full study population. We then define and outline a framework to generate synthetic negative control exposures to detect bias in LASSO PS weighted analyses.
@@ 2.4.1 Negative Control Exposures
In addition to the data structure, $\{Y,X,A\}$, defined previously, assume we observe a binary random variable, $Z$, generated with probability $g(X)=P(Z|X)$. Let $Y^{z}$ represent the potential outcome under the condition $Z=z$, and $Y^{(a,z)}$ the potential outcome that would be observed under the condition $A=a$ and $Z=z$. We define $Z$ to be a negative control exposure if the following conditions hold: 1) there is no causal relationship between the negative control exposure and the outcome; 2) there is no unmeasured common cause between the negative control exposure and the outcome; 3) there is no association between the negative control exposure and the exposure conditional on $X$, and 4) the conditional distribution of $A$ given $X$ is structurally equivalent to the conditional distribution of $Z$ given $X$ in the sense that $e(X)$ and $g(X)$ contain the same information with respect to causal relationships between $X$ and $A$ and between $X$ and $Z$. This can be stated more formally in the following definition.
@@
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Y$, $X$, and $A$ are defined in Section 2.1. Let $Z$ be a binary random variable generated with probability $g(X)=P(Z|X)$, where $10<g(X)<$ 1. We define $Z$ to be a negative control exposure if for all $a$ and $z$: 1) $Y^{(a,z)}=Y^{a}$; 2) $Y^{a}\perp\!\!\!\perp Z|X$; 3) $Z\perp\!\!\!\perp A|X$; and 4) $g(X)=f(e(X))$, where $f(\cdot)$ is any 1-to-1 function of $e(X)$, where $e(X)=P(A|X)$.
It is important to note that there are many alternative definitions for a negative control exposure [36]. Which definition is most appropriate depends on how the negative control exposure is used. Definition 1 follows similar conditions to those of a disconnected negative control defined in Kummerfeld et al [37] and Shi et al [36]. The difference being that Condition 4 in Definition 1 (structural equivalence) is written only in terms of measured covariates $X$ and assumes no unmeasured confounding. In addition, the definition we use here is more restrictive since the assumption of structural equivalence (Condition 4) requires that there is a unique mapping between $e(X)$ and $g(X)$. This ensures that $e(X)$ and $g(X)$ are balancing scores for both $A$ and $Z$ [6], which implies that the set of all covariates and higher order terms (including interactions) affecting exposure and the negative control exposure are equivalent. This condition implies, but is even stronger, than Markov Equivalence where two conditional probability distributions, $e(X)$ and $g(X)$, share the same causal graph with respect to conditional independencies between $X$ and $A$ and between $X$ and $Z$ [38, 39]. Figure 1 illustrates a directed acyclic graph for a negative control exposure that satisfies Definition 1.
$A$$Y$$Z$$X$
Figure 1: Directed acyclic graph illustrating possible causal relationships between a set of covariates ($X$), a negative control exposure ($Z$), an exposure ($A$) and an outcome ($Y$).
Definition 1 implies that if conditioning on a set of covariates is sufficient to satisfy exchangeability in the study population, then conditioning on the same set of covariates will satisfy conditional independence between the observed outcome, $Y$, and the negative control exposure, $Z$, in the study population. This is formally stated in the following proposition. The proof to Proposition 1 follows immediately from the definition of a negative control exposure (details are provided in the Supplemental Appendix).
@@
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure. If $Y^{a}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X$.
Proposition 1 implies that negative control exposures can be used for bias detection. If adjustment for a set of covariates, $X$, does not result in unbiased null effect estimates when evaluating the association between a negative control exposure and the observed outcome, then under the assumption of no unmeasured confounding (for both the exposure and negative control exposure), then the analytic approch does not sufficiently control for all confounder information in $X$ (e.g., a LASSO PS weighted analysis with improper undersmoothing). It is important to note that negative controls are typically used for bias detection of unmeasured confounding which requires additional assumptions on the causal structure of unmeasured confounders between exposure and the negative control exposure. Here, we are simply interested in the use of negative control exposures to detect bias caused by the lack of control for measured confounders under the assumption of no unmeasured confounding.
Although negative control exposures can be used for bias detection, Definition 1 makes no statement on how covariates in, $X$, relate to the negative control exposure in terms of strength and direction of effects. Therefore, it is difficult to know how the magnitude of bias caused by $X$ on the effect of $Z$ on $Y$ relates to the magnitude of bias caused by $X$ on the effect of $A$ on $Y$ without additional assumptions on how $g(X)$ relates to $e(X)$. Still, even if $g(X)=e(X)$ so that the distribution of covariates across the exposure groups match those across the negative control exposure groups, the structure and magnitude of bias caused by $X$ on the relationship between $A$ and $Y$ can differ substantially from that between $Z$ and $Y$ if $A$ is causally related to the outcome (since $Z$ has a backdoor path to the outcome through $A$, where the only backdoor path from $A$ to $Y$ is through $X$). If one conditions on $A$, the magnitude of bias caused by $X$ on the relationship between $A$ and $Y$ versus that between $Z$ and $Y$ can still differ if covariate effects on the outcome vary between strata of $A$ [40].
@@
If one wishes to avoid making assumptions on the causal relationship between $A$ and $Y$ when using negative control exposures to understand the structure and magnitude of confounding bias, one could restrict on $A=0$ so that the confounding structure between $Z$ and $Y$ among the unexposed approximates the confounding structure between $A$ and the counterfactual outcome, $Y^{a=0}$, in the study population. After restricting to the unexposed population a similar result to Proposition 1 holds. Specifically, it can be shown that conditional independence between a negative control exposure, $Z$, and the observed outcome, $Y$, within the unexposed population is connected to conditional independence between the actual exposure, $A$, and counterfactual outcome, $Y^{a=0}$, in the full study population (i.e., partial exchangeability). This can be stated more formally in the following proposition. The proof to Proposition 2 is provided in the Supplemental Appendix.
@@
Assume we have the data structure, $\{Y,X,A,Z\}$ defined previously, where $Z$ is a negative control exposure. If $Y^{a=0}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X,A=0$.
Proposition 2 implies that if adjusting for a set of variables, $X$, is sufficient to satisfy partial exchangeability in the full study population, then proper adjustment for the same set of variables will satisfy conditional independence between the observed outcome, $Y$, and the negative control exposure, $Z$, in the unexposed population. This result is similar to Proposition 1, but since we are restricting to the unexposed population, the negative control exposure can only be used to detect violations of partial exchangeability (as opposed to full conditional exchangeability).
As mentioned previously, the purpose for restricting to unexposed individuals when estimating the effect of $Z$ on $Y$, is to approximate the confounding structure between $A$ and the counterfactual outcome, $Y^{a=0}$, in the study population. Similarity between these two confounding structures depends only on how $g(X)$ relates to $e(X)$. If $Y^{a=0}$ could be observed in the full study population and $g(X)=e(X)$ so that the distribution of covariates across negative control exposure groups match those across the exposure groups, then it is straightforward that the magnitude of bias caused by $X$ on the effect of $Z$ on $Y^{a=0}$ will be equal the magnitude of bias caused by $X$ on the effect of $A$ on $Y^{a=0}$ [40].
In practice, however, $Y^{a=0}$ can only be observed for those where $A=0$. After restricting on $A=0$, it is not possible for the distribution of covariates across the negative control exposure groups to match those across the exposure groups in the full population. Still, as long as the odds of exposure are proportional to the odds of the negative control exposure, general patterns in the separation of covariates across exposure groups will be similar to those across the negative control exposure groups after restricting to the unexposed (i.e., restricting on $A=0$). We define a negative control exposure with proportional odds as follows:
Definition 2.
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Y$, $X$, and $A$ are defined in Section 2.1, and $Z$ is a binary random variable that satisfies the conditions for a negative control exposure as stated in Definition 1. We say that $Z$ is a negative control exposure with proportional odds if $\frac{g(X)}{1-g(X)}\propto\frac{e(X)}{1-e(X)}$.
Comment 1.
Definition 2 satisfies the conditions in Definition 1, but is more restrictive. Specifically, in addition to requiring a 1-to-1 mapping between $e(X)$ and $g(X)$, the proportional odds assumption places restrictions on how the strength and direction of covariate effects on exposure relate to the strength and direction of covariate effects on the negative control exposure. Specifically, if the $logit(e(X))$ follows Equation (1) (i.e. is linear in $X$), the proportional odds assumption ensures that the coefficient for each covariate (conditional effect) in the linear predictor for the $logit(e(X))$ and the $logit(g(X))$ is equivalent, with the only difference between the linear predictors potentially being the intercept.
Comment 2.
The goal of the proportional odds assumption is to have covariate differences across negative control exposure groups within the unexposed that approximate covariate differences across the exposure groups in the full population. If covariate differences across exposure groups match those across negative control exposure groups within the unexposed, then under the assumption that the association between $X$ and $A$ is linear, it can be shown that the magnitude of confounding bias caused by $X$ on the relationship between $A$ and $Y^{a=0}$, will be equal to the bias caused by $X$ on the relationship between $Z$ and the observed outcome, $Y$, among the unexposed on the absolute scale (see proof in Supplemental Appendix S1.4). If the relationship between $X$ and $A$ is not linear, but is linear in some expanded basis (e.g., the Highly Adaptive Lasso), then the same arguments hold and are generally applicable by replacing $X$ with the expanded set of basis functions.
Comment 3.
@@
In practice, however, the proportional odds assumption only guarantees that covariate differences across negative control exposure groups within the unexposed will approximate differences across exposure groups in the full study population; they will not be exact. Still, even when covariate differences across negative control exposure groups within the unexposed do not mirror those of the full exposed and unexposed populations, we conjecture that the strength of confounding caused by covariates on the relationship between $Z$ and $Y$ among the unexposed will be approximately proportional to the strength of confounding caused by covariates on the relationship between $A$ and $Y^{a=0}$ in the study population. Therefore, when using negative control exposures within the unexposed to detect bias in LASSO PS weighted estimators, we suggest evaluating both the bias and the percent bias (bias relative to the bias of the unadjusted estimate) to better understand how the bias within negative control exposure studies relates to the study cohort.
@@ 2.4.2 A framework for generating synthetic negative control exposure cohorts
Let $i=1,...,n$ index a sample of $n$ independent and identically distributed random variables with data structure, $\{Y,X,A\}$, where $Y$, $X$, $A$, and $e(X)$ are defined in Section 2.1. Let $\widehat{e}(X)_{n}$ represent a sample estimator for $e(X)$. We define a synthetic negative control exposure with proportional odds as a binary variable that is generated with probability $\widehat{g}(X)_{n}$, where $\frac{\widehat{g}(X)_{n}}{1-\widehat{g}(X)_{n}}\propto\frac{\widehat{e}(X)_{n}}{1-\widehat{e}(X)_{n}}$.
@@
Several variations for generating synthetic negative control exposures have been proposed [41, 42, 43, 44]. A common thread across the approaches is that they approximate a confounding structure where exchangeability between the synthetic exposure and observed outcome among the unexposed is closely connected to partial exchangeability in the full study population, as described in Proposition 2. Here, we propose additional modifications to previous frameworks for the purpose of validating LASSO PS-weighted analyses. We outline the framework in the following algorithm and provide comments on various aspects of the framework below.
@@
5:For each sampled unit, $j$ (where $j=1,...,n$), in the $k^{th}$ bootstrapped sample, conduct a single independent Bernoulli trial with probability $\pi_{j}$ to determine whether sampled unit $j$ is assigned to the synthetic exposure group.
@@
Comment 1.
The framework begins by fitting a LASSO model predicting exposure that is tuned using cross-validation. This model does not implement any undersmoothing and, consequently, is not expected to minimize bias in PS weighted causal estimators. However, this is not the objective since this model is not used for constructing causal estimators. The goal here is to fit the most accurate model for the PS function, $e(X)$, to generate synthetic negative control exposures that mimic the causal structure between covariates and exposure. Ertefaie et al [15] explain that while undersmoothing can result in PS weighted estimators that converge at a faster rate to causal parameters, undersmoothing does not improve the rate of convergence to the PS function, $e(X)$, itself. A LASSO model that is tuned using cross-validation will estimate the PS regression more precisely, whereas an undersmoothed LASSO model sacrifices some precision of that estimate in the interests of enhancing precision of estimation of the eventual causal parameter. Therefore, we do not undersmooth the fitted LASSO model in Step 1 to avoid degrading the accuracy of the fitted model in terms of approximating the true PS function, $e(X)$.
Comment 2.
@@ 3 Evaluation
The first simulation setup was motivated from Wyss et al [20] where the data structure was simulated to reflect settings common in healthcare database studies where the prevalence of exposure is much greater than the outcome incidence and where the vast majority of baseline features available for covariate adjustment are spurious binary indicators (sparse high-dimensional data structures). The second simulation setup comes directly from one of the simulations in Ertefaie et al [15]. This second simulation included a smaller set of baseline covariates, but where the distribution of the baseline covariates and their association with exposure were more complex (continuous covariates with nonlinear associations with exposure). We briefly outline each simulation setup below. Details along with R software code are provided in the Supplemental Appendix.
  • Simulation Setup 1: We simulated a binary exposure, $A$, a binary outcome, $Y$, and 1000 baseline covariates, where $e(X)=expit(\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\ldots+\beta_{100}X_{100})$ and $P(Y|X)=expit(\alpha_{0}+\alpha_{1}X_{1}+\alpha_{2}X_{2}+\ldots+\alpha_{100}X_{100})$. Each coefficient in the exposure and outcome model was drawn from a separate $unif(0,0.693)$ distribution. Each $X_{k}\sim Bernoulli(0.2)$, where ($k=1,\ldots,100$). We simulated an additional 900 variables binary variables ($X_{101}$ through $X_{1000}$) that had no effect on exposure or outcome (spurious variables). The prevalence of exposure and incidence of the outcome were simulated to be 30% and 5%, respectively. We considered sample sizes of 5,000, 10,000, 20,000, and 40,000.
@@
For both simulations, we generated 500 datasets. For Simulation 1, we estimated the PS using LASSO regression that included the main effects of all 1000 variables. For Simulation 2, we used the Highly Adaptive Lasso with default tuning parameters from the hal9001 R package [53] to generate an expanded set of indicator basis functions, $W$. We then fit LASSO models with different tuning criteria on the generated features, $W$. Tuning approaches included cross-validation and the previously described balance metrics. For each fitted model, we estimated the mean difference in outcomes across exposure groups using the following PS weighting approaches [54, 55]. Example R code is provided in the Supplemental Appendix.
  • Inverse Probability Weighting: $w_{i}=\frac{A_{i}}{e(X_{i})}+\frac{(1-A_{i})}{(1-e(X_{i}))}$
[…]
  • Overlap Weights: $w_{i}=A_{i}(1-e(X_{i}))+(1-A_{i})e(X_{i})$
Previous work has shown that cross-fitting or using out-of-sample predictions when estimating nuissance functions can improve the properties of effect esitmates [20, 56, 57, 58]. Therefore, for all LASSO models, we used out-of-fold (out-of-sample) predictions when assigning weights.
For each simulated dataset, we then ran Algorithm 1 described previously to produce synthetic negative control exposures for bias detection. For Step 4 in Algorithm 1, we took 500 bootstrapped samples from each simulated dataset to generate 500 synthetic negative control exposure cohorts. We applied analyses to each of the generated synthetic cohorts to calculate the estimated bias and percent bias for that simulated dataset. We then averaged across all the simulated datasets to evaluate how well the synthetic bias and synthetic percent bias aligned with the actual bias and percent bias for the two simulation setups described above.

### v3

@@ 2.4 Using Synthetic Data for Bias Detection with LASSO PS Analyses
Because the synthetic cohorts do not approximate the full data distribution, they are simply used to detect analyses that are unlikely to satisfy partial exchangeability (discussed in Section 2.4.1). Using synthetic cohorts for bias detection rather than model selection helps to mitigate the challenge of making analytic decisions that overfit to the synthetic data. In Section 2.4.1, we formally define negative control exposures and describe how they relate to conditional exchangeability in the full study population. We then outline a framework to generate synthetic negative control exposures to detect bias in LASSO PS weighted analyses.
@@ 2.4.1 Negative Control Exposures
In addition to the data structure, $\{Y,X,A\}$, defined previously, assume we observe a binary random variable, $Z$, generated with probability $g(X)=P(Z|X)$. Let $Y^{z}$ represent the potential outcome under the condition $Z=z$, and $Y^{(a,z)}$ the potential outcome that would be observed under the condition $A=a$ and $Z=z$. Here, we define $Z$ to be a negative control exposure if the following conditions hold:
  1. 1.
    There is no causal relationship between $Z$ and the outcome, $Y$.
  2. 2.
    There is no unmeasured common cause between $Z$ and $Y$.
  3. 3.
    There is no association between $Z$ and the exposure, $A$, conditional on $X$.
  4. 4.
    The odds of $A$ given $X$ are proportional to the odds of $Z$ given $X$.
These conditions are stated more formally in the following definition:
@@
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Y$, $X$, and $A$ are defined in Section 2.1. Let $Z$ be a binary random variable generated with probability $g(X)=P(Z|X)$, where $0<g(X)<1$. We define $Z$ to be a negative control exposure if for all $a$ and $z$: 1) $Y^{(a,z)}=Y^{a}$; 2) $Y^{a}\perp\!\!\!\perp Z|X$; 3) $Z\perp\!\!\!\perp A|X$; and 4) $\frac{g(X)}{1-g(X)}\propto\frac{e(X)}{1-e(X)}$.
It is important to note that there are many alternative definitions for a negative control exposure [40]. Which definition is most appropriate depends on how the negative control exposure is used. Definition 1 follows similar conditions to those of a disconnected negative control defined in Kummerfeld et al [41] and Shi et al [40], with two important differences. First, Definition 1 defines negative control exposures in terms of their relationship with measured covariates, $X$, rather than unmeasured covariates, $U$. This is because for our purposes here, we are interested in the use of negative control exposures for bias detection caused by measured confounders under the assumption of no unmeasured confounding. Second, the definition we use here is more restrictive due to Condition 4 (the proportional odds assumption). Specifically, Condition 4 requires that there is a unique mapping between $e(X)$ and $g(X)$ (1-to-1 function). This ensures that $e(X)$ and $g(X)$ are balancing scores for both $A$ and $Z$ [6], which implies that the set of all covariates (and higher order terms) affecting the exposure and negative control exposure are equivalent. The proportional odds assumption further places restrictions on how the strength and direction of covariate effects on $A$ relate to the strength and direction of covariate effects on $Z$. Specifically, if the $logit(e(X))$ follows Equation (1) (i.e. is linear in $X$), the proportional odds assumption ensures that the coefficient for each covariate (conditional effect) in the linear predictor for the $logit(e(X))$ and the $logit(g(X))$ is equivalent, with the only difference between the linear predictors potentially being the intercept. Figure 1 illustrates one example of a causal graph for a negative control exposure that satisfies Definition 1.
[…]
Figure 1: Directed acyclic graph illustrating possible causal relationships between an exposure (AA), an outcome ($Y$), a set of covariates ($X_{1}$ and $X_{2}$), and a negative control exposure ($Z$).
Definition 1 implies that if conditioning on a set of covariates is sufficient to satisfy exchangeability in the study population, then proper adjustment for the same set of covariates and exposure will satisfy conditional independence between the observed outcome, $Y$, and the negative control exposure, $Z$, in the study population. This is formally stated in the following proposition. The proof of Proposition 1 follows immediately from the definition of a negative control exposure (details are provided in the Supplemental Appendix).
@@
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure that satisfies Definition 1. For any $X_{s}\subseteq X$, if $(Y^{a=1},Y^{a=0})\perp\!\!\!\perp A|X_{s}$ then $Y\perp\!\!\!\perp Z|X_{s},A$.
Proposition 1 implies that negative control exposures can be used for detecting bias caused by measured confounders. If an analysis is used to select and adjust for a set of covariates when estimating the effect of $Z$ on $Y$, bias in the estimated effect would indicate that the analysis does not adequately control for all confounder information in $X$. Such bias could occur because: 1) the analysis does not select a sufficient adjustment set, or 2) the analysis selects a sufficient adjustment set but the analytic approach does not adequately control for the selected covariates (e.g., bias caused by improper tuning of LASSO PS models).
It is important to emphasize that when conducting analyses using negative control exposures, conditioning on $A$ is necessary to block associations between $Z$ and $Y$ that go through the exposure, $A$. Failing to adjust for $A$ can result in the structure and magnitude of bias caused by $X$ on the relationship between $Z$ and $Y$ being different from that between $A$ and $Y$, with sufficient adjustment sets not being identical. For example, in Figure 1, adjusting for $X_{1}$ is sufficient to remove confounding for the association between $A$ and $Y$, but it is not sufficient to remove confounding between $Z$ and $Y$ due to the backdoor path from $Z$ to $Y$ through $X_{2}$ and $A$. While conditioning on $A$ results in the confounding structure (and sufficient adjustment sets) between $Z$ and $Y$ being identical to that between $A$ and $Y$, the magnitude of confounding caused by each covariate can differ across these two confounding structures if covariate effects on the outcome vary between strata of $A$ [42]. Consequently, establishing a connection between the relative strength of confounding caused by each covariate on the association between $Z$ and $Y$ with that between $A$ and $Y$ requires strong assumptions on how $A$ affects $Y$ (including how $A$ modifies the effect of $X$ on $Y$).
@@
If one wishes to avoid making strong assumptions on the causal relationship between $A$ and $Y$ when using negative control exposures to understand the structure and magnitude of confounding bias, one could instead restrict on $A=0$ when estimating the effect of $Z$ on $Y$. The purpose of restricting to unexposed individuals when estimating the effect of $Z$ on $Y$, is to approximate the confounding structure between $A$ and the counterfactual outcome, $Y^{a=0}$, in the study population. Similarity between these two confounding structures depends only on how covariate distributions across exposure groups in the study population relate to those across the negative control exposure groups within the unexposed (requires no assumptions on how the exposure and covariates affect the outcome). After restricting to the unexposed population a similar result to Proposition 1 holds. Specifically, it can be shown that if adjusting for a set of covariates is sufficient to satisfy partial exchangeability in the full study population, then proper adjustment for the same set of covariates will satisfy conditional independence between the observed outcome, $Y$, and the negative control exposure, $Z$, in the unexposed population. This can be stated more formally in the following proposition. The proof to Proposition 2 is provided in the Supplemental Appendix.
@@
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure that satisfies Definition 1. For any $X_{s}\subseteq X$, if $Y^{a=0}\perp\!\!\!\perp A|X_{s}$ then $Y\perp\!\!\!\perp Z|X_{s},A=0$.
@@
Proposition 2 establishes that the covariate sets sufficient for achieving partial exchangeability for the effect of $A$ on $Y$ in the study population are also sufficient to remove confounding for the association between $Z$ and $Y$ within the unexposed group. However, the degree of confounding bias caused by covariates will not be identical in both cases. This is because covariate distributions across exposure groups cannot perfectly mirror those across negative control exposure groups after restricting on $A=0$. Still, since the proportional odds assumption ensures that the conditional effect of each covariate on the exposure and negative control exposure are equal (on the odds ratio scale), covariate differences along with the relative strength and patterns of confounding across these two scenarios will be similar (discussed further in Supplemental Appendix S1.4). Therefore, when using negative control exposures for bias detection, we suggest evaluating both the bias and the percent bias (bias relative to the bias of the unadjusted estimate) to better understand how the bias within negative control exposure studies reflects bias in the study cohort.
@@ 2.4.2 A framework for generating synthetic negative control exposure cohorts
Let $n$ represent the sample size for a set of independent and identically distributed observations with data structure, $\{Y,X,A\}$, where $Y$, $X$, AA, and $e(X)$ are defined in Section 2.1. Let $\widehat{e}(X)_{n}$ represent a sample estimator for $e(X)$. We define a synthetic negative control exposure as a binary random variable that is generated with probability $\widehat{g}(X)_{n}$, where $\frac{\widehat{g}(X)_{n}}{1-\widehat{g}(X)_{n}}\propto\frac{\widehat{e}(X)_{n}}{1-\widehat{e}(X)_{n}}$.
@@
Several variations for generating synthetic negative control exposures have been proposed [43, 44, 45, 46]. A common thread across the approaches is that they approximate a confounding structure where exchangeability between the synthetic exposure and observed outcome among the unexposed is closely connected to partial exchangeability in the full study population, as described in Proposition 2. Here, we propose additional modifications to previous frameworks for the purpose of bias detection for LASSO PS-weighted analyses. We outline the framework in Algorithm 1. Comments on various aspects of the framework are provided below.
@@
5:For each sampled unit, $j$ (where j=1,,nj=1,...,n), in the $k^{th}$ bootstrapped sample, conduct a single independent Bernoulli trial with probability $\pi_{j}$ to determine whether sampled unit $j$ is assigned to the synthetic exposure group.
@@
The framework begins by fitting a LASSO model predicting exposure that is tuned using cross-validation. This model does not implement any undersmoothing and, consequently, is not expected to minimize bias in PS weighted estimators. However, this is not the objective since this model is not used for constructing PS weighted estimators. The goal here is to fit the most accurate model for the PS function, $e(X)$, to generate synthetic negative control exposures that mimic the causal structure between covariates and exposure. Ertefaie et al [9] explain that while undersmoothing can result in PS weighted estimators that converge at a faster rate to causal parameters, undersmoothing does not improve the rate of convergence to the PS function, $e(X)$, itself. A LASSO model that is tuned using cross-validation will estimate the PS regression more precisely, whereas an undersmoothed LASSO model sacrifices some precision of that estimate in the interest of enhancing precision of estimation of the eventual causal parameter. Therefore, we do not undersmooth the fitted LASSO model in Step 1 to avoid degrading the accuracy of the fitted model in terms of approximating the true PS function, $e(X)$.
@@ 3 Evaluation
The first simulation setup was motivated from Wyss et al [11] where the data structure was simulated to reflect settings common in healthcare database studies where the prevalence of exposure is much greater than the outcome incidence and where the vast majority of baseline features available for covariate adjustment are spurious binary indicators (sparse high-dimensional data structures). The second simulation setup comes directly from one of the simulations in Ertefaie et al [9]. This second simulation included a smaller set of baseline covariates, but where the distribution of the baseline covariates and their association with exposure were more complex (continuous covariates with nonlinear associations with exposure). We briefly outline each simulation setup below. Details along with R software code [55] are available on GitHub [56].
  • Simulation Setup 1: We simulated a binary exposure, $A$, a binary outcome, $Y$, and 1000 baseline covariates, where $e(X)=expit(\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\ldots+\beta_{100}X_{100})$ and $P(Y|X)=expit(\alpha_{0}+\alpha_{1}X_{1}+\alpha_{2}X_{2}+\ldots+\alpha_{100}X_{100})$. Each coefficient in the exposure and outcome model was drawn from a separate uniform distribution in $[0,0.693]$. Each $X_{k}\sim Bernoulli(0.2)$, where ($k=1,\ldots,100$). We simulated an additional 900 variables from separate Bernoulli(0.2) distributions ($X_{101}$ through $X_{1000}$) that had no effect on exposure or outcome (spurious variables). The prevalence of exposure and incidence of the outcome were simulated to be 30% and 5%, respectively. We considered sample sizes of 5,000, 10,000, 20,000, and 40,000.
@@
For both simulations, we generated 500 datasets. For Simulation 1, we estimated the PS using LASSO regression that included the main effects of all 1000 variables. For Simulation 2, we used the Highly Adaptive LASSO with default tuning parameters from the hal9001 R package [57] to generate an expanded set of indicator basis functions, $W$. We then fit LASSO models with different tuning criteria on the generated features, $W$. Tuning approaches included cross-validation and the previously described balance metrics. For each fitted model, we estimated the mean difference in outcomes across exposure groups using the following PS weighting approaches [58, 59]:
  • Inverse Probability Weighting (IPW): $w_{i}=\frac{A_{i}}{e(X_{i})}+\frac{(1-A_{i})}{(1-e(X_{i}))}$
  • Matching Weights (MW): $w_{i}=A_{i}\left[\frac{\min(e(X_{i}),\>1-e(X_{i}))}{e(X_{i})}\right]+(1-A_{i})\left[\frac{\min(e(X_{i}),\>1-e(X_{i}))}{(1-e(X_{i}))}\right]$
  • Overlap Weights (OW): $w_{i}=A_{i}(1-e(X_{i}))+(1-A_{i})e(X_{i})$
As discussed previously, cross-fitting or using out-of-sample predictions is important to reduce problems caused by overfitting when using machine learning to construct estimators for causal effects [9, 21, 22, 23]. Therefore, for all LASSO models, we used the out-of-fold (out-of-sample) predictions when assigning weights.
For each simulated dataset, we ran Algorithm 1 described previously to produce synthetic negative control exposures for bias detection. For Step 4 in Algorithm 1, we took 500 bootstrapped samples from each simulated dataset to generate 500 synthetic negative control exposure cohorts. We applied analyses to each of the generated synthetic cohorts to calculate the estimated bias and precent bias for that simulated dataset. We then averaged across all the simulated datasets to evaluate how well the synthetic bias and synthetic percent bias aligned with the actual bias and percent bias for the two simulation setups described above. While our primary focus is on bias for reasons discussed previously (see Section 2.4.2), we also calculated the standard deviation and 95% coverage probability for each estimator to better understand how the different undersmoothing metrics impact the variability and coverage rates of PS weighted estimators.